OpenStudy (anonymous):

1+1/x / 1/x - x , x is not equall to zero please!

OpenStudy (anonymous):

\[\frac{1+\frac{1}{x}}{\frac{1}{x}-x}\] Is that right? what happens when you multiply the whole thing by X/X?

OpenStudy (anonymous):

yes thats exactly wht it looks like!

OpenStudy (anonymous):

how do i solve it

OpenStudy (anonymous):

There is no equals sign. The best you can do is simplify it... unless is it equal to something?

OpenStudy (anonymous):

ohk so its not equivalent to anything ?

OpenStudy (anonymous):

They would have to have made the thing =0 or something. You can simplify it... which is what I am assuming that they want you to do.

OpenStudy (anonymous):

ok thank you (: what is its 1/3+1/3x / 1/x + 1/3 ? what should i do?

OpenStudy (anonymous):

It's equivalent to its simplified form, tatyana, but there's no need to introduce an equals sign if it didn't originally come with one.

OpenStudy (anonymous):

are you explaining the first equation or second?

OpenStudy (anonymous):

Are there 2 equations?

OpenStudy (anonymous):

how do i solve 1/3+1/3x / 1/x + 1/3

OpenStudy (anonymous):

Ok, try to multiply by a fancy form if 1 that will get rid of all the fractions. Probably the lowest common denom. of all the fractions divided by itself. For example \[\frac{1+\frac{1}{3x}}{1+\frac{1}{5x}}\to \] \[\frac{1+\frac{1}{3x}}{1+\frac{1}{5x}}\frac{15x}{15x}\to \] \[\frac{15x+\frac{15x}{3x}}{15x+\frac{15x}{5x}}\to \] \[\frac{15x+\frac{5}{1}}{15x+\frac{3}{1}}\to \]

OpenStudy (anonymous):

ok imma try!